OpenStudy (anonymous):

Express 0.00002003701 in scientific notation

OpenStudy (anonymous):

In scientific notation, you take 2 steps to the right of the decimal point and you express it as a power of 10. 10 = 10 10^(-1) = 1 10^(-2) = 0.1 2 * 10 = 20 2 * 10^(-1) = 2 2 * 10^(-2) = 0.2 Thus: 0.00002003701 = 2.00 * 10^(-?)